I have two questions:
1. -8f < 48
2.128/-g < 4

Answer by math-vortex(648)

Hi, there--

THE PROBLEM:
Solve for f:
-8f < 48

A SOLUTION:
We solve inequalities using very similar rules to those for solving equations. One BIG difference:
When you divide both sides on an inequality by a negative number, the direction of the inequality must 
change. In this case.

Divide both sides of the inequality by -8 to isolate the variable.
-8f/-8 > 48/-8

Simplify.
f > -6

In other words, any number greater than -6 will make the inequality true.

Try f = -5.
-8f < 48
-8(-5) < 48
40 < 48
TRUE

Try f = 0.
-8f < 48
-8(0) < 48
0 < 48
TRUE.

What is we pick a number less than or equal to -6. It will not make the inequality true.
Let f = -6
-8f < 48
-8(-6) < 48
48 < 48
FALSE

Try f = -10.
-8f < 48
-8(-10) < 48
80 < 48
Also FALSE.

Hope this helps! Feel free to email if you have any questions about the solution.

Good luck with your math,

Mrs. F
math.in.the.vortex@gmail.com